Blue Angel, Inc., a private firm in the holiday gift industry, is considering a new project. The company currently has a target debt–equity ratio of .35, but the industry target debt–equity ratio is .30. The industry average beta is 1.90. The market risk premium is 7 percent, and the risk-free rate is 5 percent. Assume all companies in this industry can issue debt at the risk-free rate. The corporate tax rate is 34 percent. The project requires an initial outlay of $694,000 and is expected to result in a $114,000 cash inflow at the end of the first year. The project will be financed at the company’s target debt–equity ratio. Annual cash flows from the project will grow at a constant rate of 5 percent until the end of the fifth year and remain constant forever thereafter. Calculate the NPV of the project. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) NPV $
Answer #1

Cost of capital=0.35/1.35*5%*(1-34%)+1/1.35*(5%+7%*1.90/(1+(1-34%)*0.30)*(1+(1-34%)*0.35))=14.68248933%

NPV=-694000+114000/(1+r)+114000/(1+r)*((1+g)/(1+r))+114000/(1+r)*((1+g)/(1+r))^2+114000/(1+r)*((1+g)/(1+r))^3+114000/(1+r)*((1+g)/(1+r))^4+114000/(1+r)*((1+g)/(1+r))^4*1/(r-0%)

=-694000+114000/(1+14.68248933%)+114000/(1+14.68248933%)*((1+5%)/(1+14.68248933%))+114000/(1+14.68248933%)*((1+5%)/(1+14.68248933%))^2+114000/(1+14.68248933%)*((1+5%)/(1+14.68248933%))^3+114000/(1+14.68248933%)*((1+5%)/(1+14.68248933%))^4+114000/(1+14.68248933%)*((1+5%)/(1+14.68248933%))^4*1/(14.68248933%-0%)

=201637.7385